# Constrained Portfolio Analysis in High Dimensions: Tracking Error and Weight Constraints

### v1

@@ A.2 Relation Between Tracking Error and Risk Aversion
$TE\geq\sqrt{(\boldsymbol{w}^{*}-\boldsymbol{m})^{\prime}\boldsymbol{\Sigma}_{y}(\boldsymbol{w}^{*}-\boldsymbol{m})}=\sqrt{\boldsymbol{w}_{d}^{*\prime}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{d}^{*}}$. From equations (1)-(5) of Bajeux-Besnainou et al. (2011) we have, $\boldsymbol{\Theta}\mathrel{\mathop{:}}=\boldsymbol{\Sigma}_{y}^{-1}$:
@@
Define $\delta=\frac{{\boldsymbol{1}_{p}}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}}{{\boldsymbol{1}_{p}}^{\prime}\boldsymbol{\Theta}{\boldsymbol{1}_{p}}}$, then $\boldsymbol{w}_{d}^{*}=\frac{\boldsymbol{\Theta}}{\Xi}(\boldsymbol{\mu}-\delta{\boldsymbol{1}_{p}})$. Now we plug in $TE$ and by symmetry of $\boldsymbol{\Theta}$
$TE$ $\geq\sqrt{\boldsymbol{w}_{d}^{*\prime}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{d}^{*}}=\sqrt{\frac{1}{\Xi}\cdot(\boldsymbol{\mu}-\delta{\boldsymbol{1}_{p}})^{\prime}\boldsymbol{\Theta}\cdot\boldsymbol{\Sigma}_{y}\cdot\frac{1}{\Xi}\boldsymbol{\Theta}\cdot(\boldsymbol{\mu}-\delta{\boldsymbol{1}_{p}})}$
$=\frac{1}{\Xi}\cdot\sqrt{(\boldsymbol{\mu}-\delta{\boldsymbol{1}_{p}})^{\prime}\cdot\boldsymbol{\Theta}\cdot(\boldsymbol{\mu}-\delta{\boldsymbol{1}_{p}})}$
@@
A.3 Joint Tracking Error and Binding Weight Constraints
@@
by p.345 of Horn and Johnson (2013) for the first inequality, and we use $l_{1}$ norm of transpose of a matrix is equal to $l_{\infty}$ norm of a matrix for the first equality, and for the second inequality, we use $l_{\infty}$ norm matrix definition, and for the rate we use Theorem 2 of Caner et al. (2023), and $10<r/p\leq C_{1}<$ 1 with $C_{1}$ a positive constant below 1. Then by reverse triangle inequality
@@
By Assumption we know that $|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}/p|\geq c>0$, and
@@
Then clearly by (A.48)(A.49), and by assuming $|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}|/p\geq c>0$,
@@
Next, in the second term in (A.45) by the analysis in (A.33), and $10<r/p<$ 1
$\|\boldsymbol{\Theta}\boldsymbol{1}_{R}/p\|_{1}=O_{p}(\bar{s}r_{T}K^{3/2}).$ (A.52)
@@
$\left\|\boldsymbol{\Theta}\boldsymbol{1}_{R}/p\right\|_{1}\left|\frac{1}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}/p}-\frac{1}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}/p}\right|=O_{p}(\bar{s}l_{T})O_{p}(\bar{s}r_{T}K^{3/2}).$ (A.53)
@@
$(\hat{w}_{k}-\hat{w}_{a})-(w_{k}-w_{a})$ $=$ $\left[\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}\right]$ (A.54)
@@
$\left[\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}\right]$ $=$ $\left[\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}\right]$ (A.55)
$+$ $\left[\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}\right].$
@@
$\left|\left[\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}\right]\right|$$\leq$ $\left|\frac{(\boldsymbol{1}_{R}^{\prime}(\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta})^{\prime}\boldsymbol{1}_{R})/p}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p}\right|$ (A.56)
@@
where $\boldsymbol{M}\mathrel{\mathop{:}}p\times p$ generic matrix and $\boldsymbol{x}\mathrel{\mathop{:}}p\times 1$ generic vector, $\boldsymbol{M}_{j}$ as the jth row of $M$, (written in column form: $\boldsymbol{M}_{j}\mathrel{\mathop{:}}p\times 1$) we have, by Holder’s inequality and the inequality immediately above
@@
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{p}\right|$$\leq$ $\|\boldsymbol{1}_{R}\|_{\infty}\|\boldsymbol{\Theta}\boldsymbol{1}_{R}\|_{1}/p$ (A.59)
$\leq$ $\|\boldsymbol{\Theta}\|_{l_{1}}\|\boldsymbol{1}_{R}\|_{1}/p$
$=$ $\|\boldsymbol{\Theta}\|_{l_{\infty}}\|\boldsymbol{1}_{R}\|_{1}/p$
$=$ $\|\boldsymbol{\Theta}\|_{l_{\infty}}r/p$
$=$ $O_{p}(\bar{s}r_{T}K^{3/2}),$
where we use Holder for the first inequality, then p.345 of Horn and Johnson (2013) for the second inequality, and symmetry of the precision matrix for the first equality, and (A.10) for the rate since $10<r/p<$ 1. Then combine (A.57)(A.58)(A.59) (A.51) in first right side term in (A.54), where the second term in (A.56) has the slower rate
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}\right|=O_{p}(\bar{s}^{2}l_{T}r_{T}K^{3/2}).$ (A.60)
@@
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}\right|=O_{p}(\bar{s}^{2}l_{T}r_{T}K^{3/2}).$ (A.61)
@@
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}/p}{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}\right|=O(\bar{s}r_{T}K^{3/2}).$ (A.63)
@@
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}\right|=O(\bar{s}r_{T}K^{3/2}).$ (A.64)
@@
$\|\boldsymbol{k}-\boldsymbol{a}\|_{1}\leq\frac{\|\boldsymbol{\Theta}\boldsymbol{1}_{R}\|_{1}/p}{|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}|/p}+\frac{\|\boldsymbol{\Theta}\boldsymbol{1}_{p}\|_{1}/p}{|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}|/p}=O(\bar{s}r_{T}K^{3/2})+O(\bar{s}r_{T}K^{3/2}),$
by the analysis in (A.52) and $|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}|/p\geq c>0$ assumption.Q.E.D.
Lemma A.2.
@@
Then set for easing the derivations and proofs $\boldsymbol{y}\mathrel{\mathop{:}}=(\boldsymbol{k}-\boldsymbol{a})\mathrel{\mathop{:}}p\times 1,\hat{\boldsymbol{y}}\mathrel{\mathop{:}}=(\hat{\boldsymbol{k}}-\hat{\boldsymbol{a}})\mathrel{\mathop{:}}p\times 1,x\mathrel{\mathop{:}}=w_{k}-w_{a},\hat{x}\mathrel{\mathop{:}}=\hat{w}_{k}-\hat{w}_{a}$, where $\hat{x},x$ are both scalars. So we have
@@
Now use Lemma A.1(iii)-(iv) and by definition of $\hat{x}\mathrel{\mathop{:}}=\hat{w}_{k}-\hat{w}_{a},x\mathrel{\mathop{:}}=w_{k}-w_{a}$, and by Assumption $|w_{k}-w_{a}|\geq c>0$, for a positive constant $c$,
$|\hat{x}x|\geq c^{2}-O(\bar{s}r_{T}K^{3/2})O_{p}(\bar{s}^{2}r_{T}l_{T}K^{3/2})\geq c^{2}-o_{p}(1),$ (A.68)
by Assumption 8, $\bar{s}^{3}r_{T}^{2}l_{T}K^{3}=o(1)$. Then consider the numerator of the first right side in (A.66), by $\hat{\boldsymbol{y}},\boldsymbol{y},\hat{x},x$ definitions at the beginning of the proof
@@
$=$ $O_{p}(\bar{s}^{2}l_{T}r_{T}K^{3/2}).O_{p}(\bar{s}^{2}l_{T}r_{T}K^{3/2})=o_{p}(1),$
by Lemma A.1(i)-(iii) and triangle inequality, and Assumption 8, $\bar{s}^{2}l_{T}r_{T}K^{3/2}=o(1)$. Next, we consider the numerator of the second right side term in (A.66) by $y,\hat{x}-x$ definitions
@@
$=$ $O(\bar{s}r_{T}K^{3/2})O_{p}(\bar{s}^{2}l_{T}r_{T}K^{3/2})$
$=$ $O_{p}(\bar{s}^{3}r_{T}^{2}l_{T}K^{3})=o_{p}(1),$
where we use Lemma A.1(iii),(v) and Assumption 8. Combine (A.67)-(A.70) in (A.66) and use definitions of $\hat{y},y,\hat{x},x$ at the beginning of the proof here
@@
$=$ $O_{p}(\bar{s}^{3}r_{T}^{2}l_{T}K^{3})=o_{p}(1),$
where the rate is determined by (A.70) by Assumption 8. Consider the second term on the right side of (A.65), with $\hat{\boldsymbol{y}}-\boldsymbol{y}\mathrel{\mathop{:}}=(\hat{\boldsymbol{k}}-\hat{\boldsymbol{a}})-(\boldsymbol{k}-\boldsymbol{a})$ $\hat{x}-x\mathrel{\mathop{:}}=(\hat{w}_{k}-\hat{w}_{a})-(w_{k}-w_{a})$, and $x\mathrel{\mathop{:}}=w_{k}-w_{a}$
@@
$+$ $\|\frac{(\hat{\boldsymbol{y}}-\boldsymbol{y})(\hat{x}-x)}{\hat{x}x}\|_{1}+\|\frac{(\hat{\boldsymbol{y}}-\boldsymbol{y})x}{\hat{x}x}\|_{1}$
$=$ $[O_{p}(\bar{s}^{2}l_{T}r_{T}K^{3/2})]^{2}+O_{p}(\bar{s}^{2}l_{T}r_{T}K^{3/2})O(\bar{s}r_{T}K^{3/2})$
$=$ $O_{p}(\bar{s}^{3}l_{T}r_{T}^{2}K^{3})=o_{p}(1),$
@@
$\|\frac{\hat{\boldsymbol{k}}-\hat{\boldsymbol{a}}}{\hat{w}_{k}-\hat{w}_{a}}-\frac{\boldsymbol{k}-\boldsymbol{a}}{w_{k}-w_{a}}\|_{1}=O_{p}(\bar{s}^{3}l_{T}r_{T}^{2}K^{3})=o_{p}(1).$
@@
$w_{u}\mathrel{\mathop{:}}=\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}.$

### v2

@@ A.3 Relation Between Tracking Error and Risk Aversion
$TE\geq\sqrt{(\boldsymbol{w}^{*}-\boldsymbol{m})^{\prime}\boldsymbol{\Sigma}_{y}(\boldsymbol{w}^{*}-\boldsymbol{m})}=\sqrt{\boldsymbol{w}_{d}^{*\prime}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{d}^{*}}$. From equations (1)-(5) of Bajeux-Besnainou et al. (2011) we have, $\boldsymbol{\Theta}:=\boldsymbol{\Sigma}_{y}^{-1}$:
@@
Define $\zeta:=\frac{{\boldsymbol{1}_{p}}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}}{{\boldsymbol{1}_{p}}^{\prime}\boldsymbol{\Theta}{\boldsymbol{1}_{p}}}$, then $\boldsymbol{w}_{d}^{*}=\frac{\boldsymbol{\Theta}}{\Xi}(\boldsymbol{\mu}-\zeta{\boldsymbol{1}_{p}})$. Now we plug in TE and by symmetry of $\boldsymbol{\Theta}$
$TE$ $\geq\sqrt{\boldsymbol{w}_{d}^{*\prime}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{d}^{*}}=\sqrt{\frac{1}{\Xi}\cdot(\boldsymbol{\mu}-\zeta{\boldsymbol{1}_{p}})^{\prime}\boldsymbol{\Theta}\cdot\boldsymbol{\Sigma}_{y}\cdot\frac{1}{\Xi}\boldsymbol{\Theta}\cdot(\boldsymbol{\mu}-\zeta{\boldsymbol{1}_{p}})}$
$=\frac{1}{\Xi}\cdot\sqrt{(\boldsymbol{\mu}-\zeta{\boldsymbol{1}_{p}})^{\prime}\cdot\boldsymbol{\Theta}\cdot(\boldsymbol{\mu}-\zeta{\boldsymbol{1}_{p}})}$
@@
A.4 Proofs for Joint Tracking Error and Binding Weight Constraints
@@
by p.345 of Horn and Johnson (2013) for the first inequality, and we use $l_{1}$ norm of transpose of a matrix is equal to $l_{\infty}$ norm of a matrix for the first equality, and for the second inequality, we use $l_{\infty}$ norm matrix definition, and for the rate we use Theorem 2 of Caner et al. (2023), and $0<r/p\leq C_{1}<1$ with $C_{1}$ a positive constant below 1. Then by reverse triangle inequality
@@
By Assumption 8 we know that $|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}/p|\geq c>0$, and
@@
Then clearly by (A.51)-(A.53), and by assuming $|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}|/p\geq c>0$, Assumption 8
@@
Next, in the second term in (A.49) by the analysis in (A.11), and $0<r/p<1$
$\|\boldsymbol{\Theta}\boldsymbol{1}_{R}/p\|_{1}=O(1).$ (A.56)
@@
$\left\|\boldsymbol{\Theta}\boldsymbol{1}_{R}/p\right\|_{1}\left|\frac{1}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}/p}-\frac{1}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}/p}\right|=O_{p}(\bar{s}l_{T})O(1).$ (A.57)
@@
$(\hat{w}_{k}-\hat{w}_{a})-(w_{k}-w_{a})$ $=$ $\left[\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}\right]$ (A.58)
@@
$\left[\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}\right]$ $=$ $\left[\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}\right]$ (A.59)
$+$ $\left[\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}\right].$
@@
$\left|\left[\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}\right]\right|$ $\leq$ $\left|\frac{(\boldsymbol{1}_{R}^{\prime}(\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta})^{\prime}\boldsymbol{1}_{R})/p}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p}\right|$ (A.60)
@@
where $\boldsymbol{M}:p\times p$ generic matrix and $\boldsymbol{x}:p\times 1$ generic vector, $\boldsymbol{M}_{j}$ as the jth row of $M$, (written in column form: $\boldsymbol{M}_{j}:p\times 1$) we have, by Holder’s inequality and the inequality immediately above
@@
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{p}\right|\leq C<\infty,$ (A.63)
where we use Cauchy-Schwartz inequality with (B.22) of Caner et al. (2023). Then combine (A.61)(A.62)(A.63) (A.55) in first right side term in (A.58),
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}}{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}\right|=O_{p}(\bar{s}l_{T}).$ (A.64)
@@
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}-\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}}\right|=O_{p}(\bar{s}l_{T}).$ (A.65)
@@
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}/p}{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}\right|=O(1).$ (A.67)
@@
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}\right|=O(1).$ (A.68)
@@
$\|\boldsymbol{k}-\boldsymbol{a}\|_{1}\leq\frac{\|\boldsymbol{\Theta}\boldsymbol{1}_{R}\|_{1}/p}{|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}|/p}+\frac{\|\boldsymbol{\Theta}\boldsymbol{1}_{p}\|_{1}/p}{|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}|/p}=O(1)+O(1),$
by the analysis in (A.11) and (A.56) and $|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{R}|/p\geq c>0$ Assumption 8.Q.E.D.
Lemma A.2
@@
Then set for easing the derivations and proofs $\boldsymbol{y}:=(\boldsymbol{k}-\boldsymbol{a}):p\times 1,\hat{\boldsymbol{y}}:=(\hat{\boldsymbol{k}}-\hat{\boldsymbol{a}}):p\times 1,x:=w_{k}-w_{a},\hat{x}:=\hat{w}_{k}-\hat{w}_{a}$, where $\hat{x},x$ are both scalars. So we have
@@
Now use Lemma A.1(iii)-(iv) and by definition of $\hat{x}:=\hat{w}_{k}-\hat{w}_{a},x:=w_{k}-w_{a}$, and by Assumption 8 $|w_{k}-w_{a}|\geq c>0$, for a positive constant $c$,
$|\hat{x}x|\geq c^{2}-O(1)O_{p}(\bar{s}l_{T})\geq c^{2}-o_{p}(1),$ (A.72)
by Assumption 8, $\bar{s}l_{T}=o(1)$. Then consider the numerator of the first right side in (A.70), by $\hat{\boldsymbol{y}},\boldsymbol{y},\hat{x},x$ definitions at the beginning of the proof
@@
$=$ $O_{p}(\bar{s}l_{T}).O_{p}(\bar{s}l_{T})=o_{p}(1),$
by Lemma A.1(i)-(iii) and triangle inequality, and Assumption 8, $\bar{s}l_{T}=o(1)$. Next, we consider the numerator of the second right side term in (A.70) by $\boldsymbol{y},\hat{x}-x$ definitions
@@
$=$ $O(1)O_{p}(\bar{s}l_{T})$
$=$ $O_{p}(\bar{s}l_{T})=o_{p}(1),$
where we use Lemma A.1(iii),(v) and Assumption 8. Combine (A.71)-(A.74) in (A.70) and use definitions of $\hat{\boldsymbol{y}},\boldsymbol{y},\hat{x},x$ at the beginning of the proof here
@@
$=$ $O_{p}(\bar{s}l_{T})=o_{p}(1),$
where the rate is determined by (A.74) by Assumption 8. Consider the second term on the right side of (A.69), with $\hat{\boldsymbol{y}}-\boldsymbol{y}:=(\hat{\boldsymbol{k}}-\hat{\boldsymbol{a}})-(\boldsymbol{k}-\boldsymbol{a})$ $\hat{x}-x:=(\hat{w}_{k}-\hat{w}_{a})-(w_{k}-w_{a})$, and $x:=w_{k}-w_{a}$
@@
$\leq$ $\|\frac{(\hat{\boldsymbol{y}}-\boldsymbol{y})(\hat{x}-x)}{\hat{x}x}\|_{1}+\|\frac{(\hat{\boldsymbol{y}}-\boldsymbol{y})x}{\hat{x}x}\|_{1}$
$=$ $[O_{p}(\bar{s}l_{T})]^{2}+O_{p}(\bar{s}l_{T})O(1)$
$=$ $O_{p}(\bar{s}l_{T})=o_{p}(1),$
@@
$\|\frac{\hat{\boldsymbol{k}}-\hat{\boldsymbol{a}}}{\hat{w}_{k}-\hat{w}_{a}}-\frac{\boldsymbol{k}-\boldsymbol{a}}{w_{k}-w_{a}}\|_{1}=O_{p}(\bar{s}l_{T})=o_{p}(1).$
@@
$w_{u}:=\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}.$